OpenStudy (anonymous):

Find the radical equivalent of 27^2/3 and compute it. Please help.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello!

OpenStudy (misty1212):

the exponent is a fraction \(\frac{2}{3}\) the denominator means take the cubed root the numerator means square the result

OpenStudy (misty1212):

do you know what the cubed root of 27 is?

OpenStudy (misty1212):

if you do not, say so and i will tell you

OpenStudy (anonymous):

3

OpenStudy (misty1212):

yes and \(3^2\)?

OpenStudy (anonymous):

9

OpenStudy (misty1212):

right so that is your answer \[27^{\frac{2}{3}}=3^2=9\]

OpenStudy (anonymous):

thank you so much c:

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]